• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 00:03:04 on October 27,1998

    Entry number 3792

    keyword=shift summary

    16:00-18:00 (2.796Gev perp)
    18:00-18:35 Target NMR, Pedestal run
    18:35-20:55 2.796 Gev perp
    20:55-21:03 rf trip -->no beam to hall
    21:03-21:15 beam restored 2.796 Gev perp
    21:15 completed 2.796 |_ (30.8Mevents) began momentum change,
    target NMR
    lost comm. to hadron magnets. rebooted ioc.
    22:05-22:39 magnets settled. Raster tests, target beam warmup
    22:39-24:00 data on \\2.5804Gev